A body is projected vertically upward with speed 40m/s. The distance travelled by body in the last second of upward journey is.take g as 9.8 ..why in this question u is taken 0

Grade:12th pass

1 Answers

Hari Sundar Pyda
6 years ago
Initial velocity= u= 40m/sg= -9.8Final velocity= v= 0We know that v=u+at0= 40-9.8t9.8t=40t= 40/9.8t= 4seconds(approximately).v=u+atv= 40-9.8(3)v= 10.6Here v is u. v2-u2= 2aS0- (10.6*10.6)= 2*-9.8*S-112.36=-19.6*SS=112.36/19.6S=5.73m= distance travelled.